OpenStudy (anonymous):

Optimization. Find the Minimum Value for the Function: f(x)=x^2-ln(x^8), where x>0. @Mathematics

OpenStudy (anonymous):

f(x) = x^2 - 8ln(x) f'(x) = 2x - 8/x 2x-8/x=0 2x=8/x x^2=4 x=2

OpenStudy (anonymous):

f(2) = 2^2 - 8ln2 point is (2, 4-8ln(2))